# Deep Learning-Enabled ISAC-OTFS Pre-equalization Design for Aerial-Terrestrial Networks

### v1

@@ V-B Effect of Iteration Number
V-C Effect of Transmit Power
@@
Fig. 7 presents the optimized CRLB versus the transmit power under different schemes, where only the sensing performance is considered (i.e., $\rho_{C}=0$). From this figure, as transmit power increases, the sensing CRLB of all schemes decreases. This is because higher transmit power indicates a higher signal-to-noise ratio for echoes, leading to improved sensing accuracy. Additionally, we can see from the figure that the performance based on our CPBPE algorithm almost overlaps with the best CRLB performance achieved by the scheme under perfect sensing parameters. When compared to the scheme based on outdated CSI for pre-equalization design, there is an improvement of around $2\mathrm{W}$ in CRLB performance. This demonstrates that prediction is necessary for achieving higher accuracy in CRLB for sensing. It is worth noting that even with the scheme based on outdated sensing parameters, significant sensing performance gains compared to the scheme without pre-equalization can still be obtained. This validates the ability of the proposed pre-equalization design module to achieve sensing performance gains even under less accurate estimation.
@@
Figure 7: Sensing CRLB performance under different transmit power ($\rho_{C}=0$).
V-D Trade-off Between Communication and Sensing Performance
Fig. 8 illustrates the trade-off between communication and sensing performance achieved by different algorithms. As seen in the figure, when the communication weighting factor $\rho_{C}$ is large, the proposed algorithm achieves an MSE of the order of $10^{-3}$. As the communication weighting factor $\rho_{C}$ decreases, the sensing CRLB of the CPBPE algorithm is reduced by more than 65 %, indicating a significant improvement in sensing accuracy. This demonstrates that the designed weighting factor allows different pre-equalization designs to adapt to varying communication and sensing requirements. Moreover, the trade-off curve of the proposed CPBPE algorithm nearly overlaps with that of the design based on perfect CSI and sensing parameters, while the trade-off curve of the pre-equalization design based on outdated CSI always lies outside the trade-off curve of the proposed CPBPE algorithm. This means that, under any communication and sensing requirements, the proposed method can overcome the impact of outdated CSI, achieving a good balance between communication and sensing performance. Note that the CRLB performance of the pre-equalization design based on outdated CSI improves overall with the increase in communication symbol demodulation MSE. However, when the communication weighting factor $\rho_{C}$ is very small, the CRLB increases. This is because when the gap between outdated and perfect CSI and sensing parameters is larger, the better the optimization result based on the outdated CSI, the poorer communication and sensing performance when applied with real CSI. This again underscores the necessity of obtaining accurate CSI.
@@ V-D Trade-off Between Communication and Sensing Performance
V-E Evaluation of Signal Processing Complexity at the UE
In traditional communication frameworks as shown in Fig. 2, the UAV UE needs to perform channel equalization before symbol detection, which requires matrix inversion operations. This results in high computational complexity, i.e., $O((MN)^{3}+2^{o}MN)$, where $o$ is the modulation order. In contrast, the proposed CPBPE scheme only requires a complexity of $O(2^{o}MN)$ at the UE. The comparison of the signal processing complexity of the proposed framework with the traditional framework (i.e., performing channel equalization at the UE side) is plotted in Fig. 9. As shown in the figure, the processing complexity increases with the increase in symbol length. Moreover, the complexity gap between the traditional communication scheme and the proposed CPBPE scheme widens significantly. It is worth noting that given that energy consumption is proportional to complexity, the proposed method greatly reduces energy consumption at the UE, thereby enhancing the battery life of the UAV.
@@ VI Conclusions
In this work, we have proposed a pre-equalization design based on the OTFS waveform for an ISAC-enabled aerial-terrestrial network. To overcome the influence of outdated CSI and sensing parameters, relying on deep learning, a CPBPE framework was developed to effectively design a more precise pre-equalization matrix. The pre-equalization design enabled the receiver to require only simple power normalization for symbol detection, significantly reducing the overhead and complexity associated with channel estimation and equalization at the UE. Simulation results showed that the CPBPE framework enhances both communication and sensing performance compared to pre-equalization based on outdated CSI. Moreover, it closely approximated the performance achieved with perfect CSI and sensing parameters, demonstrating its robustness and suitability for high-mobility scenarios.

### v2

@@ V-B Convergence Behavior
V-C Impact of Transmit Power
@@
V-D Trade-off Between Communication and Sensing Performance
Fig. 7 illustrates the trade-off between communication and sensing performance achieved by different algorithms. As seen in the figure, when the communication weighting factor $\rho_{C}$ is large, the proposed algorithm achieves an MSE of the order of $10^{-2}$. As the communication weighting factor $\rho_{C}$ decreases, the sensing CRLB of the CP-PE algorithm is reduced by more than 60%, indicating a significant improvement in sensing accuracy. This demonstrates that the designed weighting factor allows different pre-equalization designs to adapt to varying communication and sensing requirements. Moreover, the trade-off curve of the proposed CP-PE algorithm nearly overlaps with that of the design based on perfect CSI and sensing parameters, while the trade-off curve of the pre-equalization design based on outdated CSI always lies outside the trade-off curve of the proposed CP-PE algorithm. This means that, under any communication and sensing requirements, the proposed method can overcome the impact of outdated CSI, achieving a good balance between communication and sensing performance. Note that the CRLB performance of the pre-equalization design based on outdated CSI improves overall with the increase in communication symbol demodulation MSE. However, when the communication weighting factor $\rho_{C}$ is very large, the MSE increases. This is because when the gap between outdated and perfect CSI and sensing parameters is larger, the better the optimization result based on the outdated CSI, the poorer communication and sensing performance when applied with real CSI. This again underscores the necessity of obtaining accurate CSI.
@@ V-D Trade-off Between Communication and Sensing Performance
V-E Communication Receiver Complexity Comparison
In traditional communication frameworks as shown in Fig. 2, the UAV UE needs to perform channel equalization before symbol detection, which requires matrix inversion operations. This results in high computational complexity, i.e., $O((MN)^{3}+2^{o}MN)$, where $o$ is the modulation order. In contrast, the proposed CP-PE framework only requires a complexity of $O(2^{o}MN)$ at the UE. The comparison of the signal processing complexity, represented by the total number of complex multiplications, between the proposed framework and the traditional framework (i.e., performing channel equalization at the UE side) is plotted in Fig. 8. As shown in the figure, the processing complexity increases with the increase in symbol length. Moreover, the complexity gap between the traditional communication scheme and the proposed CP-PE framework widens significantly. Even for short symbol lengths, the processing complexity is reduced by more than 75% compared to channel equalization at the UE side. It is worth noting that, given that energy consumption is proportional to complexity, the proposed method greatly reduces energy consumption at the UE, thereby enhancing the battery life of the UAV.
@@ VI Conclusions
In this work, we proposed a pre-equalization design based on the OTFS waveform for an ISAC-enabled air-ground network. To overcome the influence of outdated CSI and sensing parameters, relying on deep learning, a CP-PE framework was developed to effectively design a more precise pre-equalization matrix. The pre-equalization design enabled the receiver to require only simple power normalization for symbol detection, significantly reducing the overhead and complexity associated with channel estimation and equalization at the UE. Simulation results showed that the CP-PE framework enhances both communication and sensing performance compared to pre-equalization based on outdated CSI. Moreover, it closely approximated the performance achieved with perfect CSI and sensing parameters, demonstrating its robustness and suitability for high-mobility scenarios.